OpenStudy (anonymous):

Write the expression: 1. the sum of 6 and t = 6+ t a. the product of s and 12

OpenStudy (anonymous):

1. is correct....... the sum of 6 and t = 6+t 2. the product of s and 12 = s*12 = 12s

OpenStudy (anonymous):

whenever you multiply a number and a variable, the number is written first and the variable later... ex the product of m and (m*5) = m * 5m = 5m^2